OpenStudy (anonymous):

Solve \[x^5-x+13=0\] show your steps

OpenStudy (anonymous):

sorry it's \[x^5 -x -13=0\]

OpenStudy (precal):

Using a calculator it seems that there is only one real zero, x is approximately -1.712123 I would think the other 4 solutions might be complex.

OpenStudy (precal):

Ok even so the zero is now positive instead of negative.

jhonyy9 (jhonyy9):

x(x4 -1)=13 x(x2 +1)(x2 -1)=13 x(x2 +1)(x-1)(x+1)=13 x((x+1)2 -2x))(x-1)(x+1)=13

OpenStudy (precal):

I am looking at your second to the last step since I disagree with your solution (and I disagree with your last step) \[x(x ^{2}+1)(x-1)(x+1)=13\] This is incorrect because you must set the equation to zero and factor using all terms (this means including the -13).